OpenStudy (jojo12g):

Simplify. 27 ÷ (–9 – (–6 )) A. 9 B. 3 C. –3 D. –9

OpenStudy (jojo12g):

i say d

OpenStudy (anonymous):

So first tackle the part inside the brackets: (-9 - (-6)) Subtracting a negative makes a positive so: = (-9 + 6) = (6 - 9) (rearranging) 6-9 = 3 Now, 27 ÷ (-9 - (-6)) = 27 ÷ (3) = 9 A

OpenStudy (anonymous):

^^ She's right.

OpenStudy (jojo12g):

thanks